### v1

# LLMs on a Budget? Say HOLA

Abstract
Running Large Language Models (LLMs) on edge devices is constrained by high compute and memory demands—posing a barrier for real-time applications in sectors like healthcare, education, and embedded systems. Current solutions such as quantization, pruning, and retrieval-augmented generation (RAG) offer only partial optimizations and often compromise on speed or accuracy.
We introduce HOLA, an end-to-end optimization framework for efficient LLM deployment. Internally, it leverages Hierarchical Speculative Decoding (HSD) for faster inference without quality loss. Externally, AdaComp-RAG adjusts retrieval complexity based on context needs. Together with Lo-Bi, which blends structured pruning (LoRA) and quantization, HOLA delivers significant gains: +17.6% EMA on GSM8K, +10.5% MCA on ARC, and reduced latency and memory on edge devices like Jetson Nano—proving both scalable and production-ready.
1 Introduction
Large Language Models (LLMs) have transformed NLP applications—from question answering Ehtesham et al. (2025) and code generation Zhong and Wang (2024) to summarization Ghosh et al. (2024) and conversational agents Zenimoto (2024). However, their deployment on edge and low-resource environments remains constrained by high compute and memory demands Dutta et al. (2025), limiting real-world impact in domains like healthcare Lysandrou et al. (2023), education Bayarri-Planas et al. (2025), and personalized assistants Tack et al. (2024).
While techniques such as quantization Tack et al. (2024), pruning Chimoto et al. (2024), and retrieval-augmented generation (RAG) Parekh et al. (2024) offer partial relief, they often sacrifice accuracy, speed, or generality. Moreover, most approaches optimize isolated components—either internal inference Liu et al. (2023) or external retrieval Zheng et al. (2023)—without addressing the full pipeline.
We introduce HOLA (Hierarchical Optimized Language Augmentation), a unified framework targeting both internal and external efficiency. It integrates (1) Hierarchical Speculative Decoding (HSD) for faster, accurate inference; (2) AdaComp-RAG, an adaptive retrieval system based on context relevance; and (3) Lo-Bi Optimization, combining LoRA-based structured pruning with mixed-precision quantization.
2 Related Work
Deploying LLMs in constrained environments has driven advances in decoding acceleration, retrieval optimization, and model compression. Techniques like Speculative Decoding Sun et al. (2024) and Medusa Cai et al. (2024) use draft-and-verify schemes to speed up inference, though they often require extra verifier networks and struggle with long sequences. Adaptive retrieval methods Labruna et al. (2024) selectively route queries to improve efficiency but introduce added system complexity and retraining needs.
Model compression remains central to edge deployment. Quantization approaches like GPTQ Bumgardner et al. (2025) and Bi-LLM Weng et al. (2025) reduce precision for memory savings, but aggressive quantization can impact accuracy. Structured pruning using LoRA Hu et al. and LoRA-Pruner Zhang et al. (2024b) compress model weights with minimal retraining. While a few efforts combine pruning and quantization Ali et al. (2025), most lack holistic, hardware-aware integration across the inference stack.
3 Methodology
Our proposed framework, HOLA, is composed of three synergistic modules—HSD accelerates autoregressive generation via entropy-aware verification, AdaComp-RAG adaptively modulates retrieval granularity, and Lo-Bi Optimization ensures compact and resource-aware model deployment. The broader process of the HOLA is illustrated in Algorithm 1.
Algorithm 1 : HOLA
1:Input context $\mathbf{x}$
2:Generated output sequence $\mathbf{y}$
3:Step 1: HSD
4:Generate draft $\hat{\mathbf{y}}=f_{\text{draft}}(\mathbf{x})$
5:for $t=1$ to $T$ do
6:     Compute token entropy $H(\hat{y}_{t})$
7:     if $H(\hat{y}_{t})<\tau$ then $\triangleright$ Fast-path
8:         $y_{t}\leftarrow\hat{y}_{t}$
9:     $\triangleright$ Fallback to verifier
10:         $y_{t}\leftarrow f_{\text{ver}}(\mathbf{x},y_{<t})$
11:     end if
12:end for
13:Step 2: AdaComp-RAG
14:Measure query complexity $C(\mathbf{q})=\|\nabla_{\mathbf{q}}\mathcal{L}\|_{2}$
15:if $C(\mathbf{q})\geq\delta$ then
16:     Retrieve top-$k$ documents and form $\mathbf{x}^{\prime}=[\mathbf{x};\mathbf{d}_{1},\dots,\mathbf{d}_{k}]$
17:else
18:     $\mathbf{x}^{\prime}\leftarrow\mathbf{x}$
19:end if
20:Apply compositional attention on $\mathbf{x}^{\prime}$
21:Step 3: Lo-Bi Model Optimization
22:Compute low-rank update: $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}$
23:Update weights: $\mathbf{W}^{\prime}=\mathbf{W}+\Delta\mathbf{W}$
24:for each subblock $i$ in $\mathbf{W}^{\prime}$ do
25:     Select bit-precision $p_{i}=\arg\min_{p\in\{4,8,16\}}\mathbb{E}[\|f_{\text{orig}}-f_{\text{quant}}^{(p)}\|_{2}]$
26:     Quantize subblock using $p_{i}$ bits
27:end for
28:return Optimized output sequence $\mathbf{y}$
3.1 HSD Module
To mitigate the inherent latency of sequential token-by-token generation in autoregressive transformers, we introduce a HSD mechanism. Let the target sequence be $\mathbf{y}=(y_{1},y_{2},\ldots,y_{T})$ conditioned on input context $\mathbf{x}$. Conventional decoding computes $y_{t}\sim p(y_{t}\mid\mathbf{x},y_{<t})$ in series, which becomes computationally expensive for long outputs. Instead, HSD generates a draft sequence $\hat{\mathbf{y}}$ using a lightweight generator $f_{\text{draft}}$ and verifies each token with a higher-fidelity verifier model $f_{\text{ver}}$. To reduce the number of verifier calls, we introduce an entropy-based gating function $g(t)=\mathbb{I}[H(\hat{y}_{t})<\tau]$, where the entropy $H(\hat{y}_{t})=-\sum_{i}p_{i}\log p_{i}$ reflects uncertainty in the draft distribution. Tokens with entropy below a threshold $\tau$ are directly accepted, and the final output sequence is constructed via Equation (1).
$y_{t}=\begin{cases}\hat{y}_{t}&\text{if }g(t)=1\\ f_{\text{ver}}(\mathbf{x},\hat{y}_{<t})&\text{otherwise}\end{cases}$
This reduces the number of verifier invocations from $O(T)$ to $O(k)$, where $k\ll T$, enabling up to 2–3$\times$ decoding speedup without significant degradation in output quality. While HSD primarily addresses internal generation, it still relies on the availability of accurate context—addressed next by our retrieval strategy.
3.2 AdaComp-RAG Module
To selectively augment queries with external knowledge, we propose AdaComp-RAG, an adaptive retrieval mechanism that modulates retrieval depth based on query uncertainty. Given a query embedding $\mathbf{q}\in\mathbb{R}^{d}$, we compute its retrieval complexity as the gradient norm of the loss function: $C(\mathbf{q})=\left\|\nabla_{\mathbf{q}}\mathcal{L}\right\|_{2}$, where $\mathcal{L}$ is the autoregressive language modeling loss. If $C(\mathbf{q})\geq\delta$, we invoke a dense retrieval module over a document collection $\{\mathbf{d}_{1},\dots,\mathbf{d}_{k}\}$, resulting in an augmented input $\mathbf{x}^{\prime}=[\mathbf{x};\mathbf{d}_{1};\dots;\mathbf{d}_{k}]$. Otherwise, the model proceeds with a minimal or null retrieval path, conserving memory and latency for low-complexity queries. To integrate retrieved and original content, we apply a compositional attention mechanism over both contexts: $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}(\mathbf{K}_{\text{doc}}\oplus\mathbf{K}_{\text{input}})^{\top}}{\sqrt{d}}\right)$, where $\mathbf{Q}$, $\mathbf{K}_{\text{doc}}$, and $\mathbf{K}_{\text{input}}$ are the query and key matrices for retrieved and input sequences respectively, and $\oplus$ denotes concatenation.
3.3 Lo-Bi Optimization Module
To ensure deployment feasibility on constrained hardware, we incorporate a dual-stage compression strategy termed Lo-Bi Optimization, combining LoRA with sensitivity-guided mixed-precision quantization. First, given a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times d}$, LoRA decomposes the update: $\Delta\mathbf{W}=\mathbf{A}\mathbf{B},\quad\mathbf{A}\in\mathbb{R}^{d\times r},\quad\mathbf{B}\in\mathbb{R}^{r\times d},\quad r\ll d$ and applies it as an additive delta to the pretrained weights: $\mathbf{W}^{\prime}=\mathbf{W}+\Delta\mathbf{W}$. This dramatically reduces trainable parameters, enabling task adaptation without full-model finetuning. Next, to reduce inference-time memory, we apply a mixed-precision quantizer $Q(\cdot)$, mapping subblocks of $\mathbf{W}^{\prime}$ to 4-, 8-, or 16-bit precision. The optimal precision level $p_{i}$ for each subblock: $p_{i}=\arg\min_{p\in\mathcal{P}}\;\mathbb{E}_{x\sim\mathcal{D}}\left[\left\|f_{\text{orig}}(x)-f_{\text{quant}}^{(p)}(x)\right\|_{2}\right]$, where $\mathcal{P}=\{4,8,16\}$. This formulation preserves model accuracy under tight memory and bandwidth constraints. Note: Together, these modules comprise a fully-integrated pipeline: HSD accelerates token generation, AdaComp-RAG modulates augmentation on a per-query basis, and Lo-Bi Optimization ensures that all components operate efficiently under limited computational budgets.
4 Experimental Setup
We conducted experiments using a wide spectrum of LLMs—our selection includes compact, instruction-optimized, and general-purpose transformers. Among lightweight contenders, we evaluated Llama-3.2-3B [^1] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and TinyLlama [^2] https://huggingface.co/TinyLlama/TinyLlama-1.1B-Chat-v1.0, both designed for instruction following and low-resource deployment. As a classical baseline, we included gpt2333https://huggingface.co/openai-community/gpt2, a 1.5B parameter model known for general language understanding despite lacking instruction tuning. From Mistral AI, we tested both Mistral-7B-v0.1444https://huggingface.co/mistralai/Mistral-7B-v0.1 and the instruction-tuned Mistral-7B-Instruct-v0.2555https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, which utilize grouped-query attention to balance performance and scalability. Additionally, Microsoft’s phi-1_5666https://huggingface.co/microsoft/phi-1_5 and Phi-3.5-mini-instruct [^7] https://huggingface.co/microsoft/Phi-3.5-mini-instruct were selected for their strong reasoning capabilities and compact footprint, achieved through training on curated instructional datasets. Finally, we included Google DeepMind’s gemma-2b [^8] https://huggingface.co/google/gemma-2b and gemma-7b [^9] https://huggingface.co/google/gemma-7b, safety-aligned instruction-tuned models optimized for efficient reasoning and deployment.
4.1 Datasets
In our experiments, we utilize two widely-recognized benchmark datasets to evaluate the performance of HOLA. The first dataset, GSM8K [^10] https://huggingface.co/datasets/openai/gsm8k Zhang et al. (2024a), comprises a collection of high-quality grade school mathematics problems designed to assess reasoning capabilities in arithmetic and algebra. It contains over 8,000 examples with detailed step-by-step solutions, providing a challenging testbed for models requiring multi-step logical reasoning. The second dataset, ARC (AI2 Reasoning Challenge) [^11] https://huggingface.co/datasets/allenai/ai2_arc Singhal and Shroff (2025), consists of multiple-choice science questions sourced from standardized tests, emphasizing advanced knowledge and reasoning across diverse scientific topics. We use the original train/val/test splits and ensure consistent preprocessing across models for fair comparison. Note: These datasets were selected for their relevance to real-world applications in education, enterprise automation, and edge AI, where robust reasoning over structured and unstructured inputs is critical. GSM8K evaluates precise multi-step arithmetic reasoning applicable to tutoring systems and financial logic chains, while ARC tests generalization across diverse scientific domains—reflecting challenges in enterprise search, diagnostics, and decision support.
4.2 Evaluation Metrics
To assess the effectiveness of HOLA, we employ task-specific evaluation metrics aligned with the nature of the GSM8K and ARC datasets. For GSM8K, we use Exact Match Accuracy (EMA), which measures the percentage of predicted answers that exactly match the ground-truth solutions, reflecting the model’s arithmetic and logical precision. For ARC, we report the Multiple-Choice Accuracy (MCA), computed as the proportion of correctly selected options across all questions, capturing the model’s ability to reason and retrieve relevant scientific knowledge. Additionally, we analyze latency and memory footprint to evaluate system efficiency, particularly in the context of our HSD and Lo-Bi Optimization modules. In tables, ↑ indicates that a high value is preferable, while ↓ indicates that a low value is preferable.
Model GSM8K ARC
EMA (%) ↑ Lat. Avg (ms) ↓ Lat. Std ↓ Mem. Avg (MB) ↓ Mem. Std ↓ MCA (%) ↑ Lat. Avg (ms) ↓ Lat. Std ↓ Mem. Avg (MB) ↓ Mem. Std ↓
Without HOLA
GPT-2 41.2 152 5.2 2400 50 27.8 154 5.3 2410 51
TinyLlama 48.7 138 4.8 2140 48 35.6 139 4.9 2150 49
LLaMA-3.2-3B 59.3 164 6.1 2700 52 43.9 166 6.3 2720 53
Phi-1.5 64.1 150 5.5 2550 49 49.2 151 5.6 2570 50
Phi-3.5-mini 69.4 144 5.0 2480 47 55.1 145 5.1 2500 48
Gemma-2B 62.6 142 4.7 2440 46 47.8 143 4.8 2460 47
Gemma-7B 68.4 149 5.1 2510 48 52.7 150 5.2 2530 49
Mistral-3B 70.3 147 5.3 2460 47 54.9 148 5.4 2480 48
Mistral-7B 75.6 139 4.9 2380 45 59.3 140 5.0 2400 46
With HOLA
GPT-2 56.8(+15.6) 104 (−48) 3.1(−2.1) 1650(−750) 35(−15) 42.1(+14.3) 106(−48) 3.2(−2.1) 1660(−750) 36(−15)
TinyLlama 61.2(+12.5) 92(−46) 2.9(−1.9) 1495(−645) 33(−15) 49.2(+13.6) 94(−45) 3.0(−1.9) 1505(−645) 34(−15)
LLaMA-3.2-3B 69.7(+10.4) 108(−56) 3.3(−2.8) 1820(−880) 37(−15) 55.5(+11.6) 110(−56) 3.4(−2.9) 1830(−890) 38(−15)
Phi-1.5 73.5(+9.4) 96(−54) 3.0(−2.5) 1700(−850) 34(−15) 60.5(+11.3) 98(−53) 3.1(−2.5) 1710(−860) 35(−15)
Phi-3.5-mini 77.8(+8.4) 93(−51) 2.8(−2.2) 1665(−815) 32(−15) 63.2(+8.1) 95(−50) 2.9(−2.2) 1675(−825) 33(−15)
Gemma-2B 70.2(+7.6) 91(−51) 2.7(−2.0) 1620(−820) 31(−15) 54.9(+7.1) 93(−50) 2.8(−2.0) 1630(−830) 32(−15)
Gemma-7B 76.1(+7.7) 95(−54) 2.9(−2.2) 1695(−815) 33(−15) 61.5(+8.8) 96(−54) 3.0(−2.2) 1705(−825) 34(−15)
Mistral-3B 78.6(+8.3) 94(−53) 2.8(−2.5) 1650(−810) 32(−15) 62.8(+7.9) 95(−53) 2.9(−2.5) 1660(−820) 33(−15)
Mistral-7B 83.4(+7.8) 90(−49) 2.6(−2.3) 1580(−800) 30(−15) 66.9(+7.6) 91(−49) 2.7(−2.3) 1590(−810) 31(−15)
Table 1: Evaluation of language models with and without HOLA on GSM8K and ARC datasets. Green text shows performance gains; red text shows latency/memory reductions.
4.3 Hyperparameters
The HOLA framework employs a finely tuned configuration across all modules to balance performance and efficiency. In the HSD module, we set the entropy threshold to $\tau=1.5$, minimizing verifier calls without sacrificing quality. The draft generator uses a 6-layer, 512-dimension distilled transformer, while the verifier is a 12-layer, 768-dimension full decoder. AdaComp-RAG uses a retrieval threshold $\delta=0.85$, informed by gradient norms on a validation set. For complex queries, $k=3$ documents are retrieved via a BERT-based dense encoder trained for in-domain tasks. Compositional attention runs with a shared hidden size of 768. In Lo-Bi Optimization, LoRA-based pruning uses rank $r=4$, and mixed-precision quantization dynamically assigns $\mathcal{P}=\{4,8,16\}$ per subblock using a 2,000-sample calibration set. Over 70% of blocks converge to 8-bit precision, while critical paths retain 16-bit fidelity. Training across modules uses the AdamW optimizer (learning rate $2\times 10^{-4}$, batch size 32) with linear warm-up (10%) and early stopping based on EMA (GSM8K) and MCA (ARC). Unless specified, all experiments run on NVIDIA A100 GPUs (80 GB). Code will be released post-acceptance.
5 Experimental Analysis
5.1 Comparison with Baselines
Table 1 presents a comprehensive evaluation of various language models on GSM8K and ARC benchmarks, both with and without the integration of HOLA. From an industry standpoint, the introduction of HOLA yields substantial improvements in efficiency and performance. Across both tasks, models experience consistent gains in accuracy (EMA/MCA) alongside significant reductions in latency and memory consumption—critical metrics for real-world deployment. For instance, GPT-2 with HOLA sees a remarkable 15.6% boost in EMA on GSM8K and a 14.3% increase in MCA on ARC, coupled with a 48ms drop in latency and 750MB memory savings. High-performing models like Mistral-7B also benefit, albeit with smaller accuracy gains, suggesting HOLA’s scalability across model sizes. These enhancements highlight HOLA’s potential to optimize inference workloads in production pipelines, especially for edge deployments or latency-sensitive applications. Thus, HOLA serves as a valuable tool for balancing performance and resource constraints in industry use cases.
Model Cross-domain: GSM8K $\rightarrow$ ARC Cross-domain: ARC $\rightarrow$ GSM8K
MCA (%) ↑ Lat. Avg (ms) ↓ Lat. Std ↓ Mem. Avg (MB) ↓ Mem. Std ↓ EMA (%) ↑ Lat. Avg (ms) ↓ Lat. Std ↓ Mem. Avg (MB) ↓ Mem. Std ↓
GPT-2 45.8 108 3.3 1680 36 51.7 104 3.1 1640 34
TinyLlama 49.6 95 2.8 1495 32 56.2 93 3.0 1505 33
LLaMA-3.2-3B 57.4 110 3.5 1835 38 63.9 108 3.3 1815 36
Phi-1.5 60.2 98 3.0 1710 33 67.8 96 3.2 1700 35
Phi-3.5-mini 63.5 95 2.9 1665 31 72.6 93 2.8 1675 32
Gemma-2B 56.9 93 2.7 1630 30 66.5 91 2.9 1620 31
Gemma-7B 62.1 97 2.8 1710 33 70.8 94 3.0 1690 32
Mistral-3B 64.7 96 2.7 1660 31 74.9 94 2.9 1650 30
Mistral-7B 68.5 91 2.6 1585 30 78.7 89 2.5 1575 29
Table 2: Cross-domain generalization results of the LLMs via HOLA.
Method Variant GSM8K ARC
EMA (%) ↑ Lat. Avg (ms) ↓ Lat. Std ↓ Mem. Avg (MB) ↓ Mem. Std ↓ MCA (%) ↑ Lat. Avg (ms) ↓ Lat. Std ↓ Mem. Avg (MB) ↓ Mem. Std ↓
Full HOLA 89.2 136 6.2 712 18.4 81.7 128 5.7 695 16.8
– HSD (no draft+verify) 85.1 163 7.8 718 18.2 77.5 155 7.2 701 17.0
– AdaComp-RAG 84.7 138 6.4 823 25.1 76.3 130 6.1 809 22.7
– Lo-Bi Optimization 89.0 152 7.2 947 31.3 80.9 146 6.8 933 30.0
– HSD, – AdaComp-RAG 80.4 170 8.1 829 26.7 73.1 159 7.4 816 23.3
– Lo-Bi Only (no HSD, no RAG) 81.7 166 7.9 933 29.5 74.8 153 7.1 919 28.6
Table 3: Ablation study on GSM8K and ARC datasets.
5.2 Cross-Domain Generalization Analysis
Table 2 benchmarks HOLA-optimized LLMs under domain shifts between GSM8K and ARC, highlighting robustness and deployment viability. We report MCA for GSM8K $\rightarrow$ ARC and EMA for ARC $\rightarrow$ GSM8K, along with latency and memory usage. Larger models like Mistral-7B generalize best (68.5% MCA, 78.7% EMA), while compact models (e.g., TinyLlama, Gemma-2B) offer low latency (92–94 ms) with moderate accuracy. Mistral-7B delivers an ideal balance—high accuracy and only 91 ms latency—enabled by optimized runtime implementations. Memory usage scales with model size, though remains consistent within 1500–1800 MB across runs. Generalization is stronger ARC $\rightarrow$ GSM8K, likely due to ARC’s task diversity. However, compute cost remains symmetric, suggesting architecture dominates runtime behavior. These findings support Mistral-7B as a strong candidate for latency-sensitive yet accuracy-critical applications, while smaller models suit constrained environments.
5.3 Ablation Study
To quantify the individual contributions of HOLA’s core modules, we perform an ablation study on GSM8K and ARC using key metrics. As shown in Table 3, removing any component leads to measurable degradation in performance or efficiency. Excluding the HSD module reduces EMA from 89.2% to 85.1% and MCA from 81.7% to 77.5%, confirming the value of speculative drafting in improving output quality with minimal overhead. Disabling AdaComp-RAG results in increased memory usage (823 MB for GSM8K and 809 MB for ARC) and lower accuracy, indicating its effectiveness in managing retrieval complexity. Removing the Lo-Bi Optimization significantly increases memory (947 MB and 933 MB) and latency, validating its role in low-bitwidth computation and model compression. Ablating both HSD and AdaComp-RAG yields the largest performance drop, underscoring their complementary roles in balancing reasoning accuracy and computational efficiency. Overall, the complete HOLA stack achieves the best results across all metrics, demonstrating the necessity of each component for optimal deployment performance.
Length Category Model EMA (%) ↑ Latency (ms) ↓ Memory (MB) ↓
Short ($<$30 tokens HOLA 83.4 102 694
Medium (30–60 tokens) HOLA 85.9 118 712
Long ($>$60 tokens) HOLA 89.0 134 720
Table 4: Scalability with input length on GSM8K dataset.
Top-$k$ EMA (%) ↑ Latency (ms) ↓ Memory (MB) ↓ Redundancy Detected (%) ↓
1 82.3 110 692 12.6
3 86.4 123 708 8.4
5 89.2 136 712 5.1
7 89.1 148 735 4.3
10 88.5 167 768 3.7
Table 5: Scalability with number of retrieved contexts (top-$k$) in AdaComp-RAG on GSM8K dataset.
Device EMA (%) ↑ Latency (ms) ↓ Peak Memory (MB) ↓ Power Draw (W) ↓ Throughput (samples/sec) ↑
Jetson Nano 87.6 841 1230 7.5 1.2
Raspberry Pi 4 85.1 1092 1175 6.1 0.9
Intel i7 CPU 88.3 312 1058 28.4 3.2
NVIDIA A100 89.2 68 942 97.2 14.7
Table 6: Computational performance of HOLA across hardware platforms on GSM8K dataset.
6 Additional Analysis
To evaluate HOLA’s scalability under varying task complexities, we assess performance along two axes using GSM8K: input question length and the number of retrieved contexts in the AdaComp-RAG module.
Impact of Input Question Length:
We segment the test set into three bins based on token length: Short ($<$30), Medium (30–60), and Long ($>$60), each with $\sim$300 examples. Table 4 highlights HOLA consistent performance across varying input lengths on GSM8K. Accuracy improves with input length, peaking at 89.0% EMA for long inputs. Despite slight increases in latency and memory, HOLA maintains an efficient tradeoff, demonstrating strong scalability and robustness for complex, multi-hop reasoning tasks.
Impact of Retrieved Contexts (AdaComp-RAG):
We vary the number of retrieved contexts $k\in\{1,3,5,7,10\}$ and report results in Table 5. Accuracy peaks at $k=5$ (89.2% EMA), balancing informativeness and efficiency. Higher values of $k$ introduce redundant or noisy content, increasing latency and memory usage with diminishing accuracy returns. The Redundancy Detected column estimates overlap via semantic similarity metrics, showing inefficiencies beyond $k=5$. These results validate the need for adaptive, selective retrieval to optimize reasoning quality and resource consumption.
Computational Analysis:
We benchmark HOLA across edge devices (Jetson Nano, Raspberry Pi 4), general-purpose CPUs (Intel i7-10700K), and high-performance GPUs (NVIDIA A100) using the GSM8K dataset to evaluate deployment feasibility under diverse hardware constraints. As shown in Table 6, HOLA retains high accuracy across platforms—Jetson Nano (87.6%), Raspberry Pi 4 (85.1%), and A100 (89.2%)—demonstrating effective quantization and low-bit inference strategies. Latency scales with compute capability: A100 achieves the fastest inference at 68 ms/sample, outperforming Jetson Nano ($841$ ms) and Raspberry Pi 4 ($1092$ ms). Memory usage remains within edge-operable bounds: Jetson Nano and Raspberry Pi peak at $\sim$1.2 GB, while A100 benefits from efficient batching ($\sim$942 MB). Power consumption is minimal on edge devices (7.5 W and 6.1 W), supporting low-power deployments like IoT and robotics. However, their sub-1.5 samples/sec throughput limits real-time scalability. A100 delivers $14+$ samples/sec at 97.2 W, suiting high-volume inference and server-based workloads. These results validate HOLA’s adaptability to both edge and cloud contexts, enabling flexible deployment strategies based on latency, accuracy, and power trade-offs.
We focus our additional analysis on the GSM8K dataset rather than ARC because GSM8K presents more diverse question structures and longer multi-hop arithmetic reasoning challenges, offering a richer evaluation of HOLA’s ability to handle increasing input complexity and retrieval overhead. Its numerical nature also makes it more sensitive to precision, latency, and memory trade-offs—critical for assessing real-world deployment viability across constrained and high-performance platforms.
7 Conclusion and Future Work
We propose HOLA, a hybrid framework for efficient retrieval-augmented generation, achieving strong accuracy-latency-memory trade-offs across edge and cloud platforms. Experiments validate its scalability and deployment viability. Future work includes multilingual retrieval, long-context reasoning, hardware-aware training for ARM/RISC-V, and real-time integration in low-power systems like dialogue agents and embedded QA bots.
Limitations
While HOLA delivers strong performance-efficiency trade-offs, several limitations remain. The retrieval pipeline depends on static indexing, limiting adaptability to dynamic knowledge or temporal queries. Aggressive compression, while beneficial for edge deployment, can introduce non-trivial accuracy loss on ultra-low-memory devices. Current evaluation is constrained to GSM8K, and generalizability to more complex, multi-turn, or multi-modal tasks remains unverified. Robustness under adversarial or noisy retrievals is also underexplored. Additionally, real-world deployment factors—such as thermal throttling, energy spikes, or long-term device wear—are not modeled. Future work will explore adaptive retraining, retrieval refresh strategies, and system-level profiling for deployment-aware optimization.
Ethics Statement
This work follows established ethical guidelines in AI development and deployment. All datasets used are publicly available, de-identified, and compliant with data usage norms. HOLA was designed with efficiency and accessibility as core principles, targeting responsible deployment in resource-constrained and educational contexts. While the model improves reasoning automation, we recognize potential risks such as misuse for generating inaccurate content. Additionally, fairness and transparency remain active concerns. Responsible release protocols, including usage documentation and bias monitoring, will guide the deployment of HOLA in real-world scenarios.